OpenStudy (anonymous):

1. A number cube is rolled 24 times and lands on '2' four times and on '6' three times. A.Find the experimental probability of landing on a 2. B.Find the experimental probability of NOT landing on a 6. C. Compare the experimental probability that you found in part A to its theoretical probability. D. Compare the experimental probability that you found in part B to its theoretical probability.

OpenStudy (unklerhaukus):

A. 4 out of 24 , is 4/24 = 1/6 , one in six

OpenStudy (anonymous):

A. 1/24 b. 23/24 c. idk d. idk well nevermind. unkle probably knows more than me -_- go with his answer. once agaaaiin... xD

OpenStudy (unklerhaukus):

B. (24-3)/24 = 21/24 = 7/8

OpenStudy (unklerhaukus):

for C. and D. Assume the numbers one to six are each on a single side of the cube and the cube is fair, C. the experimental probability is equal to the theoretical probability (1/6). D. the experimental probability is 7/8= 42/48 the theoretical probability is 5/6=40/48 there were more sixes than expected

OpenStudy (anonymous):

Thanks !(: